OpenStudy (anonymous):

Find the slope of the line passing through (5, 5) and (-4, 5)

OpenStudy (anonymous):

\[m=\Delta y/\Delta x\]

OpenStudy (anonymous):

\[\left( y_2-y_1 \right)/\left( x_2-x_1 \right)\]

OpenStudy (anonymous):

\[m = \frac{y_1 - y_2}{x_1 - x_2}\]

OpenStudy (anonymous):

\[M= \frac{y2-y1}{x2-x1}\] M= the gradient/slope and so \[M = \frac{5-5}{-4-5}\] \[M=\frac{0}{-9}\] M=0

OpenStudy (anonymous):

\[5-5/-4-5\]=0/-9=0

OpenStudy (anonymous):

(5,5) and (-4,5) x1,y1 x2,y2